OpenStudy (anonymous):

Solve 3x + 2y = -12 for y.

OpenStudy (jhannybean):

Why are you spamming questions without understanding how to solve them...

OpenStudy (jhannybean):

follow the same format. http://openstudy.com/study#/updates/51c382d5e4b055e613b88003

OpenStudy (anonymous):

a) y= 3 over 2 x-6 b) y=-3 over 2 x-6 c) y=3 over 2 x-12 d) y= -3 over 2 x -12

OpenStudy (anonymous):

sorry im about to fail algebra 2 again and im just trying to get the answers so i can pass the class! but when you do explain it i actually get how to do it.. thank you!

OpenStudy (anonymous):

3x + 2y = -12 Goal: isolate the y Step 1: Subtract 3x from both sides 3x + 2y = -12 -3x -3x ----- ------ WHAT DO YOU GET ?

OpenStudy (anonymous):

@danny562 are you there??

OpenStudy (anonymous):

Well @danny562 no one can help you if you don't interact also! I have a feeling you only want the answer, since u said u can't pass the class; but let me tell you that by just getting the answers, you WONT LEARN ANYTHING! So give it a try :)

OpenStudy (anonymous):

2y=9x?